# Hidden Markov Model Filtering with Equal Exit Probabilities

### v1

@@ II-B Fixed Point Value and Learning Performance
In this section, we analyze the value of the fixed point. Since we have already proven that the reference dynamical system converges to the unique fixed point as described in Theorem 1, the position of the fixed point will be sufficient to characterize the resulting learning performance. Due to the nonlinearity of (25), explicitly solving for the exact fixed point value is challenging. Nevertheless, we can provide an estimation of its value. In Theorem 2, we derive the lower bound for the counterpart of the agent’s belief $\boldsymbol{\mu}_{i}(\cdot)$ in the reference dynamical system.
@@
for all $m=1,\ldots,M-1$ and $i=0,1,\ldots$. Here $\hat{\mu}_{i}(\theta)$ can be seen as the counterpart of agent’s belief on state $\theta$ in the reference dynamical system, and we have the following result:
@@
This result follows directly from Theorem 1 and the properties of limits.
The value of $\hat{\mu}_{0}^{\infty}$, which represents the steady-state belief on the true state in the reference dynamical system, reflects the learning performance of the algorithm: the closer $\hat{\mu}_{0}^{\infty}$ is to 1, the better the algorithm distinguishes between the correct and incorrect hypotheses at steady state; conversely, the closer $\hat{\mu}_{0}^{\infty}$ is to $1/M$, the more ambiguous the agent’s learning outcome. We provide upper and lower bound estimates for $\hat{\mu}_{0}^{\infty}$ in the following lemma.
Lemma 1
When $0<\alpha<1/M$ and the underlying true state remains unchanged, the $\hat{\mu}_{0}^{\infty}$ defined in (46) satisfies the following lower bound:
$\hat{\mu}_{0}^{\infty}\geq\frac{(1-\alpha M)\exp(\underline{d})-1+\alpha}{(1-\alpha M)(\exp(\underline{d})-1)},$ (47)
@@
$\underline{d}=\underset{m=1,\ldots,M-1}{\min}d_{m}.$ (48)
Proof:
Proof omitted due to space limitations. ∎
Remark 5
We can observe from Lemma 2 that as $\alpha\to 0$ or $\underline{d}\to\infty$, the steady-state belief on the correct state in the reference dynamical system approaches one, indicating improved learning performance. This observation aligns with intuition: better learning performance is expected in less volatile environments, with lower observation noise, and a more accurate local likelihood model. Combining the result in Lemma 2 with the fixed point values in (25) and (26), we can conclude that as $\alpha\to 0$ or $\underline{d}\to\infty$, $\hat{\mu}_{0}^{\infty}\to 1$ and the fixed point value $\hat{x}_{m}^{\infty}\to-\infty$, which means that the agent becomes more confident in trusting the underlying true state at steady state.
III From Deterministic to Stochastic Systems
In this section, we discuss the relationship between the stochastic dynamical system induced by the $\alpha$-HMM, represented by (10), and our reference deterministic dynamical system (20). In Lemma 2, we show that the expected distance between $\boldsymbol{x}_{i}$ and the fixed point of the reference dynamical system can be asymptotically bounded. Then, in Theorem 2, we compute the instantaneous error probability of the $\alpha$-HMM based on the fixed point of the reference dynamical system. To ensure the feasibility of theoretical analysis, we introduce the following assumption.
@@ III From Deterministic to Stochastic Systems
To compare the reference dynamical system and the original stochastic dynamical system, in the following we analyze two recursions:
1. The evolution of the log-belief ratio in $\alpha$-HMM, $\boldsymbol{x}_{m,i}$ ($m=1,\ldots,M-1$):
$\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})+\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (51)
2. The evolution of the counterpart of $\boldsymbol{x}_{m,i}$ in reference dynamical system, denoted by $\hat{x}_{m,i}$:
$\hat{x}_{m,i}=F_{m}(\hat{x}_{1,i-1},\ldots,\hat{x}_{M-1,i-1})-d_{m},$ (52)
where it is established in Theorem 1 that $\hat{x}_{m,i}\to\hat{x}^{\infty}_{m}$, as $i\to\infty$.
The following lemma establishes the relationship between the two dynamical systems described above:
Lemma 2
@@ 1. The evolution of the log-belief ratio in $\alpha$-HMM, $\boldsymbol{x}_{m,i}$ ($m=1,\ldots,M-1$):
$\mathbb{E}\left[\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\right]\leq O(\lambda_{1}^{i})+\frac{C}{1-\lambda_{1}},$ (53)
where $\hat{x}^{\infty}$ is as defined in (24),(25) and (26),
$\lambda_{1}=1-\min\left\{\alpha,\frac{\underline{d}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+C}\right\},$ (54)
with $\underline{d}$ and $C$ defined in (30) and (49) respectively.
Proof:
Proof omitted due to space limitations. ∎
In Lemma 2, we show that the distance between the original dynamical system and the fixed point of the reference dynamical system can be bounded. The coefficient $\lambda_{1}$ is always less than 1, then as the time $i\to\infty$, the upper bound of the expected distance between the two is related to the first absolute moment of the log-likelihood ratio of the observations and the contraction coefficient $\lambda_{1}$. In other words, when the observation noise decreases, the stochastic dynamical system corresponding to the $\alpha$-HMM (10) approaches the fixed point of the reference dynamical system.
To evaluate the learning performance of $\alpha$-HMM, we introduce the instantaneous error probability:
@@
From this definition, it is clear that the instantaneous error probability quantifies the probability of the agent failing to correctly identify the underlying true state at a given time $i$ during the online filtering process. We will then provide an estimate of this metric in the steady-state scenario for the proposed $\alpha$-HMM algorithm.
Theorem 2
When $0<\alpha<1/M$ and the underlying true state remains constant, i.e. $\theta_{i}^{\star}=\theta_{0}$ for all $i=1,2,\ldots$, the instantaneous error probability for the $\alpha-$HMM algorithm (6) satisfies:
$p_{i}^{e}\leq\frac{O(\lambda_{1}^{i})+C}{-(1-\lambda_{1})\overline{x}^{\infty}},$ (56)
@@
other variable definitions remain unchanged from those presented in Lemma 2.
Proof:
Proof omitted due to space limitations. ∎
Remark 6
@@
The contraction coefficient $\lambda_{1}$ in Theorem 3 is always less than 1, consequently the upper bound of the instantaneous error probability decays exponentially with time and eventually converges to a fixed value
$\frac{C}{-(1-\lambda_{1})\overline{x}^{\infty}}.$ (58)
@@
This result indicates that under the $\alpha$-HMM framework, the probability of erroneous learning cannot asymptotically approach zero, even in steady-state conditions, which contrasts with the behavior predicted by Bayes’ formula [7]. However, by sacrificing some learning performance, the $\alpha$-HMM framework significantly enhances adaptability, as previously discussed. The steady-state error learning probability can be reduced either by (1) pushing the equilibrium point $\hat{x}^{\infty}$ of the deterministic system further away from zero and accelerating the convergence rate (i.e., reducing $\lambda_{1}$), noting that, as shown in earlier analyses, improving the identifiability parameter $\underline{d}$ facilitates this outcome, or by (2) reducing the observation noise, thereby increasing the informativeness of the observations.
IV Numerical Experiments
IV-A Performance of $\alpha-$HMM
@@
In the first numerical experiment, we consider a filtering task where the agent observes noisy signals following a standard normal distribution. The number of possible states is set to $M=5$, and the observations follow $\boldsymbol{\xi}_{i}\sim\mathcal{N}(\theta_{i}^{\star},\sigma^{2})$. The agent’s local likelihood function model is given by $L(\boldsymbol{\xi}|\theta_{m})=\mathcal{N}(\theta_{m},\sigma^{2})$, where $\theta_{m}=m+1$, $m=0,\ldots,M-1$. We assume that every ten iterations, the underlying true state, $\theta_{i}^{\star}$, will be randomly reselected from the $M$ possible states. Under this setting, the evolution of the true state does not strictly follow a Markov chain. However, for the filtering problem in such a dynamic environment, our algorithm can still be employed to track the underlying true state.
Figure 2: (a) The likelihood model of the agent. (b) The underlying true state in each iteration. (c-e) Evolution of the belief on the true state for both $\alpha$-HMM and linearized $\alpha$-HMM under different values of $\alpha$ with $\sigma=0.5$ fixed. The success or failure of identification at different iterations is shown, i.e., whether the true state is assigned the highest belief. The black line represents the results for $\alpha$-HMM, the blue line represents the results for linearized $\alpha$-HMM, and the green numbers indicate the success rate of learning over 100 iterations. (f) Top: Learning rates for correctly identifying the true state of the two algorithms under various values of $\alpha$, evaluated via Monte Carlo simulation over 10,000 iterations. Bottom: Evolution of the belief on the true state for both algorithms under their optimal $\alpha$ settings.
@@
Fig. 2 compares the performance of the $\alpha$-HMM algorithm proposed in this paper with the algorithm employing linear log-belief ratio dynamics (18):
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\boldsymbol{\mu}^{1-\alpha M}_{i-1}(\theta_{m})L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}\boldsymbol{\mu}^{1-\alpha M}_{i-1}(\theta_{n})L(\boldsymbol{\xi}_{i}|\theta_{n})},$ (59)
referred to as the linearized $\alpha$-HMM algorithm. It is noteworthy that the belief update rule in (59) exhibits similarities to the one presented in [26], as elaborated in Remark 1. From Fig. 2, it can be observed that under various parameter settings, the accuracy of the $\alpha$-HMM algorithm is no worse than that of the linearized algorithm, and its performance with respect to $\alpha$ is more stable. A more pronounced advantage is observed when $\alpha$ is small. These results highlight the benefits of the nonlinearity inherent in the evolution of the log-belief ratio in $\alpha$-HMM algorithm.
@@
Focusing solely on the $\alpha$-HMM algorithm, when $\alpha$ is relatively small (Fig. 2(c)), the algorithm struggles to track transitions between two similar underlying true states (e.g., at iteration 40, when the true state transitions from $\theta_{2}$ to $\theta_{3}$, the algorithm fails to detect this change). However, as $\alpha$ increases, as shown in Fig. 2(d), the belief on the current true state decreases, which aligns with the analysis results in Lemma 1, allowing the algorithm to better track transitions between similar states, though with some delay. When $\alpha$ is further increased, as illustrated in Fig. 2(e), the algorithm quickly adapts to state transitions, demonstrating excellent adaptability. Nevertheless, due to the reduced confidence in the current true state, the algorithm becomes more sensitive to noise, leading to intermittent correct learning even during stable periods of the true state.

### v2

@@ 2.2 Fixed Point Value and Learning Performance
In this section, we quantify the learning performance of the $\alpha\beta$-HMM and examine its dependence on the exit probability $\alpha$, the step-size parameter $\beta$, and the identifiability parameter $\underline{d}$. Since we have already proven that the reference dynamical system converges to the unique fixed point as described in Lemma 2.2, the position of the fixed point will be sufficient to characterize the resulting learning performance. Due to the nonlinearity of (37), explicitly solving for the exact fixed point value is challenging. Nevertheless, we can provide an estimate of its value. In Theorem 2.11, we derive upper and lower bounds for the counterpart of the belief $\boldsymbol{\mu}_{i}(\cdot)$ in the reference dynamical system.
@@
for all $m=1,\ldots,M-1$ and $i=0,1,\ldots$. Here $\hat{\mu}_{i}(\theta)$ can be seen as the counterpart of belief on state $\theta$ in the reference dynamical system. By Theorem 1 and basic limit properties,
@@
The value of $\hat{\mu}_{0}^{\infty}$, which represents the steady-state belief on the true state in reference dynamical system, reflects the learning performance of the algorithm: the closer $\hat{\mu}_{0}^{\infty}$ is to 1, the better the distinction between correct and incorrect states at steady state; conversely, the closer $\hat{\mu}_{0}^{\infty}$ is to $1/M$, the more ambiguous the learning outcome. We provide upper and lower bound estimates for $\hat{\mu}_{0}^{\infty}$ in the following theorem.
Theorem 2.11.
When $0<\alpha<1/M$, $\beta>0$, and the underlying true state remains unchanged, the $\hat{\mu}_{0}^{\infty}$ defined in (73) satisfies the following bounds:
$\hat{\mu}_{0}^{\infty}\geq\frac{(1-\alpha M)\exp(\beta\underline{d})-1+\alpha}{(1-\alpha M)(\exp(\beta\underline{d})-1)}\triangleq\underline{\mu},$ (74)
$\hat{\mu}_{0}^{\infty}\leq\frac{(1-\alpha M-\alpha+\alpha/\underline{\mu})\exp(\beta\overline{d})-1+\alpha}{(1-\alpha M)(\exp(\beta\overline{d})-1)}\triangleq\overline{\mu},$ (75)
@@
$\underline{d}=\underset{m=1,\ldots,M-1}{\min}d_{m},$ (76)
and
$\overline{d}=\underset{m=1,\ldots,M-1}{\max}d_{m}.$ (77)
Proof 2.12.
Appendix 9
Remark 2.13.
The upper and lower bounds in Theorem 2.11, as functions of the parameters $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$, are illustrated in Figure 3. We observe that when $\alpha$ is smaller and $\beta\underline{d}$ is larger, the steady-state belief on the correct state in the reference dynamical system is higher, indicating improved learning performance. This observation aligns with intuition: better learning performance is expected in less volatile environments, with lower observation noise, and a more accurate local likelihood model. The step-size parameter $\beta$ and the identifiability parameter $\underline{d}$ play a decisive role; when either $\beta$ or $\underline{d}$ is large, good learning performance is ensured in deterministic dynamical system, noise-free by construction, even when $\alpha$ is relatively large. This means that, in the deterministic setting, accurate inference is possible in volatile environments, provided that the quality of the data model and observations is high enough or that the value of $\beta$ is large enough. However, choosing a large $\beta$ comes at the cost of amplifying the effect of observation noise, as we will further demonstrate in Section 3.
Figure 3: The visualizations of the upper and lower bounds given in Theorem 2.11 as functions of the parameters $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$. Top: Visualization of the lower bound function. The left and right subplots show how the lower bound varies with one parameter while the other is fixed, with $\alpha$ and $\beta\underline{d}$ set to specific values in each case. Bottom: Visualization of the upper bound function. Since the upper bound is a function of $\alpha$, $\beta\underline{d}$, and $\beta\overline{d}$ (where $\overline{d}>\underline{d}$), the three subplots illustrate how the upper bound varies with $\alpha$ and $\beta\overline{d}$ for different fixed values of $\beta\underline{d}$.
3 From Deterministic to Stochastic Systems
In this section, we discuss the relationship between the stochastic dynamical system induced by $\alpha\beta$-HMM, represented by (19), and the reference dynamical system (27), establishing that the dynamics of (27) are representative of the stochastic system (19), which describes the actual learning dynamics of the $\alpha\beta$-HMM recursion. In Lemma 3.14, we show that the expected distance between $\boldsymbol{x}_{i}$ and the fixed point of the reference dynamical system can be asymptotically bounded. Then, in Theorem 3.16, we compute the instantaneous error probability of the $\alpha\beta$-HMM based on the fixed point of the reference dynamical system. To ensure tractability of the theoretical analysis, we introduce the following assumption.
@@ 3 From Deterministic to Stochastic Systems
To compare the reference dynamical system and the original stochastic dynamical system, we recall the two recursions previously introduced:
  1. 1.
    The evolution of the log-belief ratio in $\alpha\beta$-HMM, $\boldsymbol{x}_{m,i}$ ($m=1,\ldots,M-1$):
    $\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})+\beta\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (80)
  2. 2.
    The evolution of the counterpart of $\boldsymbol{x}_{m,i}$ in the reference dynamical system, $\hat{x}_{m,i}$:
    $\hat{x}_{m,i}=F_{m}(\hat{x}_{1,i-1},\ldots,\hat{x}_{M-1,i-1})-\beta d_{m},$ (81)
    where it is established in Theorem 2.4 that $\hat{x}_{m,i}\to\hat{x}^{\infty}_{m}$, as $i\to\infty$.
The following lemma establishes the relationship between the two dynamical systems (80) and (81):
Lemma 3.14.
@@
$\mathbb{E}\left[\left\|\boldsymbol{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}\right]\leq O(\lambda_{1}^{i})+\frac{\beta C}{1-\lambda_{1}},$ (82)
where $\hat{x}^{\infty}$ is the fixed point of reference dynamical system (27),
$\lambda_{1}=1-\min\left\{\frac{2\alpha}{1-\alpha M+2\alpha},\frac{\beta\underline{d}}{2\log\frac{1-\alpha M+\alpha}{\alpha}+\beta C}\right\},$ (83)
with $C$ and $\underline{d}$ defined in (78) and (45), respectively.
Proof 3.15.
Appendix 10
In Lemma 3.14, we show that the distance between the original dynamical system and the fixed point of the reference dynamical system can be bounded. The coefficient $\lambda_{1}$ is always less than 1, then as the time instant approaches infinity, the upper bound of the expected distance between the two is related to the first absolute moment of the log-likelihood ratio of the observational model and the contraction coefficient $\lambda_{1}$. In other words, when the observation noise decreases, the stochastic dynamical system corresponding to the $\alpha\beta$-HMM (80) approaches the fixed point of the reference dynamical system.
To evaluate the learning performance of $\alpha\beta$-HMM, we introduce an important metric: the instantaneous error probability:
@@
From this definition, it is clear that the instantaneous error probability quantifies the probability of failing to correctly identify the underlying true state at a given time $i$ during the online filtering process. We will then provide an estimate of this metric in the steady-state scenario for the proposed $\alpha\beta$-HMM algorithm.
Theorem 3.16.
When $0<\alpha<1/M$, $\beta>0$, and the underlying true state remains constant, i.e. $\theta_{i}^{\star}=\theta_{0}$ for all $i=1,2,\ldots$, the instantaneous error probability for the $\alpha\beta$-HMM algorithm (5) satisfies:
$p_{i}^{e}\leq\frac{O(\lambda_{1}^{i})+\beta C}{-(1-\lambda_{1})\overline{x}^{\infty}},$ (85)
@@
other variable definitions remain unchanged from those presented in Lemma 3.14. Furthermore, we have the following asymptotic estimate of the instantaneous error probability:
If $\beta=O(\alpha)$, then:
$\lim_{\alpha\rightarrow 0}\left(\underset{i\to\infty}{\lim\sup}\;p_{i}^{e}\right)=0.$ (87)
Proof 3.17.
Appendix 11
Remark 3.18.
@@
The contraction coefficient $\lambda_{1}$ in Theorem 3.16 is always less than 1, consequently the upper bound of the instantaneous error probability decays exponentially with time and eventually converges to a fixed value
$\frac{\beta C}{-(1-\lambda_{1})\overline{x}^{\infty}}.$ (88)
@@
This result indicates that under the $\alpha\beta$-HMM framework, the probability of erroneous learning cannot be guaranteed to asymptotically approach zero for all choices of $\alpha$ and $\beta$, even in steady-state conditions, which contrasts with the behavior predicted by Bayes’ formula [8]. However, by sacrificing some learning performance, the $\alpha\beta$-HMM framework significantly enhances adaptability, as previously discussed. The steady-state error learning probability can be reduced either by (1) pushing the fixed point $\hat{x}^{\infty}$ of the deterministic system further away from zero and accelerating the convergence rate (i.e., reducing $\lambda_{1}$), noting that, as shown in earlier analyses, improving the identifiability parameter $\underline{d}$ facilitates this outcome, or by (2) reducing the observation noise, thereby increasing the informativeness of the observations.
The asymptotic analysis results for small $\alpha$ further indicate that as $\alpha$ approaches zero, the instantaneous error probability diminishes over time. This phenomenon highlights the trade-off between the learning capability and the adaptation ability of the algorithm: while a small $\alpha$ yields good learning performance, it reduces the adaptation rate, as illustrated in (83). Moreover, when $\alpha=0$, the proposed algorithm degenerates into Bayes’ formula, which, as discussed in Remark 2.9, exhibits poor adaptability.
4 Parameter Selection and Tuning
Based on the analysis in Section 2, we summarize the impacts of the exit probability parameter $\alpha$ and the step-size parameter $\beta$ as follows:
  1. 1.
    The exit probability parameter $\alpha$ can be employed to adjust the algorithm’s adaptability and learning performance. In more volatile environments, a larger value of $\alpha$ is preferred to enhance adaptability (Theorem 2.4). Conversely, if the goal is to ensure superior learning performance, a smaller value of $\alpha$ should be selected (Theorems 2.11 and 3.16). This conclusion is consistent with intuition, since $\alpha$ reflects the volatility of the environment.
  2. 2.
    The step-size parameter $\beta$ serves to balance the algorithm’s adaptability, learning performance, and the influence of observation noise. Specifically, under the assumption of the identifiability of the underlying true state (Assumption 2), a larger $\beta$ results in a faster adaptation rate (Theorem 2.4, Corollary 2.7) and improved learning performance in noise-free scenarios (Theorem 2.11). However, since $\beta$ also amplifies the effect of noise (Theorem 3.16), selecting a large $\beta$ is beneficial when the observation noise is minimal, while in high-noise conditions a large $\beta$ may lead to instability in the algorithm’s performance. Notably, when $\beta$ is set to 1, the proposed $\alpha\beta$-HMM algorithm can be interpreted as the HMM model corresponding to the equal exit probability transition matrix.
5 Numerical Experiments
5.1 Algorithm performance and parameter tuning
@@
In the first numerical experiment, we consider a filtering task where the sensor provides noisy signals following a standard normal distribution. The number of possible states is set to $M=5$, and the observations follow $\boldsymbol{\xi}_{i}\sim\mathcal{N}(\theta_{i}^{\star},\sigma^{2})$. The likelihood function model is given by $L(\boldsymbol{\xi}|\theta_{m})=\mathcal{N}(\theta_{m},\sigma^{2})$, where $\theta_{m}=m+1$, $m=0,\ldots,M-1$. We assume that every ten iterations, the underlying true state, $\theta_{i}^{\star}$, will be randomly reselected from the $M$ possible states. Under this setting, the evolution of the true state does not strictly follow a Markov chain. As we show, the $\alpha\beta$-HMM algorithm nevertheless continues to yield meaningful performance.
@@
Figure 4 compares the performance of the $\alpha\beta$-HMM algorithm with the algorithm employing linear log-belief ratio dynamics:
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\boldsymbol{\mu}^{1-\alpha M}_{i-1}(\theta_{m})L^{\beta}(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}\boldsymbol{\mu}^{1-\alpha M}_{i-1}(\theta_{n})L^{\beta}(\boldsymbol{\xi}_{i}|\theta_{n})},$ (89)
@@
referred to as the linearized $\alpha\beta$-HMM algorithm, under varying values of the parameter $\alpha$, while keeping $\beta=1$ in both algorithms to ensure a consistent influence of the data. It is noteworthy that the belief update rule in (89) exhibits similarities to the one presented in [28], as elaborated in Remark 1. From Figure 4, it can be observed that under various parameter settings, the accuracy of the $\alpha\beta$-HMM algorithm is no worse than that of the linearized algorithm, and its performance with respect to $\alpha$ is more stable. A more pronounced advantage is observed when $\alpha$ is small. These results highlight the benefits of the nonlinearity inherent in the $\alpha\beta$-HMM algorithm. Focusing solely on the $\alpha\beta$-HMM algorithm, when $\alpha$ is relatively small (Fig. 4(c)), the algorithm struggles to track transitions between two similar underlying true states (e.g., at iteration 40, when the true state transitions from $\theta_{2}$ to $\theta_{3}$, the algorithm fails to detect this change). However, as $\alpha$ increases, as shown in Fig. 4(d), the belief in the current true state decreases, allowing the algorithm to better track transitions between similar states, though with some delay. When $\alpha$ is further increased, as illustrated in Fig. 4(e), the algorithm quickly adapts to state transitions, demonstrating excellent adaptability. Nevertheless, due to the reduced confidence in the current true state, the algorithm becomes more sensitive to noise, leading to intermittent correct learning even during stable periods of the true state.
Figure 4: (a) The likelihood model. (b) The underlying true state in each iteration. (c-e) Evolution of the belief on the true state for both $\alpha\beta$-HMM and linearized $\alpha\beta$-HMM under different values of $\alpha$ with $\beta=1$ and $\sigma=0.5$ fixed. The success or failure of identification at different iterations is shown, i.e., whether the true state is assigned the highest belief. The black line represents the results for $\alpha\beta$-HMM, the blue line represents the results for linearized $\alpha\beta$-HMM, and the green numbers indicate the success rate of learning over 100 iterations. (f) Top: Learning rates for correctly identifying the true state of the two algorithms under various values of $\alpha$, evaluated via Monte Carlo simulation over 10,000 iterations. Bottom: Evolution of the belief on the true state for both algorithms under their optimal $\alpha$ settings.
The influence of the step-size parameter $\beta$ under different levels of observation noise is illustrated in Fig. 5. In the case of less noisy observations ($\sigma^{2}=1$), a higher value of $\beta$ leads to improved final learning accuracy across all three parameter settings. In contrast, when the observation noise is larger ($\sigma^{2}=2$), the learning accuracy decreases as $\beta$ increases. These experimental results are consistent with the theoretical findings presented in Theorem 3.16.
Figure 5: Likelihood models for noise variances $\sigma^{2}=1$ and $\sigma^{2}=0.5$, and the evolution of the belief on true state and the correct learning rate under different step-size parameter $\beta$ settings.
5.2 Trade-off between learning and adaption
The above results can be explained using the steady-state analysis results from Theorem 2.4 and Theorem 2.11. Assuming that $\theta_{i}^{\star}$ remains unchanged as $\theta_{0}$, i.e., $\theta_{i}^{\star}\equiv\theta^{\star}$, it is straightforward to compute:
$d_{m}=D_{\text{KL}}(L(\cdot|\theta^{\star})|L(\cdot|\theta_{m}))=\frac{(l-m)^{2}}{2\sigma^{2}},\quad\theta^{\star}=\theta_{l}.$ (90)
Thus, when $\theta^{\star}$ and $\sigma^{2}$ are fixed, the value of $d_{m}$ remains constant. In Fig. 6(a), we observe that increasing the value of $\alpha$ while keeping $\beta$ constant causes the fixed points of the reference dynamical system to approach zero, indicating lower confidence and higher sensitivity. At the same time, the convergence rate to the fixed point increases as $\alpha$ approaches $1/M$. Conversely, when $\alpha$ is fixed and $\beta$ is adjusted, larger values of $\beta$ lead to fixed points further from zero and faster convergence rates. These findings are consistent with the analytical results presented in Theorems 2.4 and 2.11.
Figure 6: (a) Comparison of the log-belief ratio evolution in the reference dynamical system under different parameter settings. Before iteration $i=25$, the underlying true state is $\theta_{i}^{=}\theta_{0}$, and after that, $\theta_{i}^{=}\theta_{4}$. The fixed point, $\hat{x}^{\infty}$, is computed using (37) when $\theta_{i}^{\star}=\theta_{0}$. (b) Comparison of the true evolution of $\left\|\hat{x}_{i}-\hat{x}^{\infty}\right\|_{\infty}$ and the evolution estimated using the contraction coefficient $\lambda$ in Theorem 2.4 on semilog y-axis coordinates. (c) Evolution of belief on the underlying true state in the reference dynamical system under different parameter settings. The estimated upper and lower bounds are computed using Theorem 2.11.
In Fig. 6(b), we further compare the convergence rates of $\hat{x}_{m,i}$ to the fixed point under different parameter settings when $\theta_{i}^{\star}\equiv\theta_{0}$. The linear decay observed on the semilog Y-axis, the upper bound estimation, and the variation in convergence rate with respect to adjustments in parameters $\alpha$ and $\beta$ are all consistent with the conclusions of Theorem 2.4 and Corollary 2.7, which predict exponential convergence to the fixed point with a rate that increases with $\alpha$ and $\beta$. In Fig.6(c), we demonstrate the evolution of the belief on the underlying true state $\theta_{0}$ in the reference dynamical system, i.e., $\hat{\mu}_{i}(\theta_{0})$, under different parameter settings. It can be observed that as $\alpha$ increases and $\beta$ decreases, both the steady-state belief on the true state and its estimated lower bound gradually decline. Furthermore, as $\alpha$ and $\beta$ increase, the belief converges to the steady state more rapidly. These results align with the analysis in Theorem 2.4 and Theorem 2.11, emphasizing the trade-off between learning performance and adaptation capability through the tuning of $\alpha$, and highlighting the role of the step-size parameter $\beta$ in shaping the convergence behavior.
5.3 Error Probability of the algorithm
In this numerical experiment, we aim to validate the conclusion from Theorem 3.16 regarding the error probability of $\alpha\beta$-HMM. We assume there are three possible states, $\Theta=\left\{\theta_{0},\theta_{1},\theta_{2}\right\}$, with the underlying true state remaining constant as $\theta_{0}$. The observations $\boldsymbol{\xi}_{i}$ are assumed to follow a truncated normal distribution with a mean of 0, specifically $\boldsymbol{\xi}_{i}=\sigma\boldsymbol{W}_{i}$, where $\sigma$ is a positive constant and $\boldsymbol{W}_{i}$ is truncated zero-mean Gaussian noise. The corresponding likelihood model under state $\theta_{m}$ assumes the observations are given by $L(\cdot|\theta_{m})=m+1+\sigma\boldsymbol{W}_{i}$. All random variables are truncated within the range $\left[-5,5\right]$.
Fig. 7(a) shows the likelihood function $L(\xi|\theta_{m})$ for all $\theta_{m}\in\Theta$ and the true distribution function of the observations $f(\xi|\theta^{\star})$ under different $\sigma$, clearly the likelihood model of state $\theta_{0}$ is closer to the true distribution. Note that there is no exact true state according to the likelihood model, as $f$ is not equal to the likelihood function on any states.